# GoLLIE : Annotation Guidelines improve Zero-Shot Information-Extraction

### 21 Feb 2024 v4

@@ 5.3 Ablation study
We have performed an ablation to see the contribution of several components in the zero-shot evaluation. We analyzed the different regularization techniques proposed in Section 3.3. Additionally, we represent the baseline, i.e when removing all components including guidelines, as ”w/o all”.
@@
The class order shuffling, guideline paraphrasing and class name masking seems to have no significant contribution to the final result, while class dropout although significant the improvements are small. As further explained in Appendix D, the loss is only computed over the result tokens, inherently limiting the model’s potential to overfit to the guidelines. In contrast, the representative annotation items give some stronger signal to the model. We see how definitions and representative candidates from the guidelines are complementary and help to improve each other.
@@ 6 Error analysis
The CrossNER dataset introduces two labels for person names within each domain. For example, in the Science domain, the labels Scientist and Person are used. The former, is used to label any person that is not a Scientist. Similarly, the Literature domain includes the labels Writer and Person. The guidelines assist GoLLIE in correctly labeling entities as Writer. However, GoLLIE still categorizes individuals as Person even when they are Scientist, despite the guidelines. This is not technically incorrect, as every scientist is, by definition, also a person.
@@
In its Political domain set, CrossNER includes the label Political Party. GoLLIE outperforms the baseline, once again demonstrating the utility of providing the model with guidelines. However, we often find that the model categorizes political parties as organizations. As listed in Table 1, most of the pretraining datasets originate from the news domain, where political parties are a common entity. However, none of the fine-tuning datasets include the Political Party entity; they are instead categorized as Organization. Consequently, during inference, the model consistently labels political parties as organizations. We believe this issue can be resolved by expanding the number and diversity of the fine-tuning datasets.
In summary, we anticipate that GoLLIE will perform well on labels with well-defined and clearly bounded guidelines. On the other hand, ambiguous labels or very coarse labels pose challenges. To this regard, be believe that GoLLIE would benefit from learning to follow instructions such as ”Label always the most specific class” or ”Annotate this class in the absence of other specific class”. We also expect that GoLLIE would benefit from expanding the number and diversity of the pre-training datasets.
@@ Acknowledgments
This work has been partially supported the Basque Government (Research group funding IT-1805-22). We are also thankful to several MCIN/AEI/10.13039/501100011033 projects: (i) DeepKnowledge (PID2021-127777OB-C21) and ERDF A way of making Europe; (ii) Disargue (TED2021-130810B-C21) and European Union NextGenerationEU/PRTR; (iii) AWARE (TED2021-131617B-I00) and European Union NextGenerationEU/PRTR. This work has also been partially funded by the LUMINOUS project (HORIZON-CL4-2023-HUMAN-01-21-101135724). Oscar Sainz is supported by a doctoral grant from the Basque Government (PRE_2023_2_0137). Rodrigo Agerri currently holds the RYC-2017-23647 fellowship (MCIN/AEI/10.13039/501100011033 and by ESF Investing in your future).
@@ Appendix A Examples
In addition to NER, EE and EAE for which examples are shown in Figures 2 and 3 respectively, we also feed the model with data from RE and SF. The formulation of RE is similar to the NER but with two argument attributes. However, the SF task is more complex as shown in the Figure 6. With this task, we added several layers of complexity to the input: extended definitions for each possible attribute (slot), optional arguments, and, fine-grained definitions of types such as Names, Values or Strings. We also added constraints into the prompt to condition the model to just output the information of our desired query instead of every single template on the text. In the future, we would like to add more complex tasks into the train and evaluation to improve the capabilities and flexibility of the model. For more examples refer to the GitHub repository.
@@ A.1 Example of generalization to new custom Tasks
Our model allows the user to define custom annotation schemas using Python code. We provide an example where we define two new types of entities: Launcher and Mission. As shown in Figure 5, Launcher and Mission are not simple entities, they correspond to what we call Template, a class similar to Entity but with additional arguments, like the SF task. For example, the space_company or the crew of the launcher are some of the additional arguments we added to the schema. As shown in the example, the model’s output (everything after result = [) satisfies the type constraints defined in the guidelines, attributes defined as strings are filled with strings and, the arguments defined as lists (like crew) are filled with lists. The model is able to correctly analyze the given sentence with our newly created annotation schema.
@@ D.1 Loss calculation
We have used the standard Next Token Prediction (NTP) loss to train our models. However, several regularizations that we applied to the models made the loss computed over the guideline tokens much higher than the actual output tokens. This is because we randomly shuffle the guidelines order, mask names or, drop classes, which makes impossible to predict what goes next. To avoid the loss of the guideline tokens overshadow the actual output tokens loss, we decided to only compute the loss over the output tokens. This way, we can also avoid some overfitting on the guidelines. This resulted on a faster training and better results overall.
@@ D.2 Dataset Detailts
Table 9 shows the number of examples for each training and zero-shot dataset. OntoNotes is generated semi-automatically and is orders of magnitude larger than the other ones. Therefore, each training epoch, we sample 30.000 random examples from the training set. The models were trained for 3 epochs with an effective batch-size of 32 and a learning-rate of 3e-4 with cosine scheduler. Therefore, we perform 15,485 training steps.
Regarding the splits, we use the standard train, dev test splits for every dataset. In case of ACE, we follow the split provided by Lin et al. (2020). In the case of CASIE we took the first 200 instances as validation and the last 2000 as test.
@@ D.4 LoRa vs Full Model Fine-Tuning
We conducted preliminary experiments to compare the performance of QLoRA (Hu et al., 2022; Dettmers et al., 2023) with that of training all the parameters in the model. These preliminary experiments were conducted using the LLaMA2 7B model Touvron et al. (2023b) and an early version of the code. However, the experimental setup for both models was identical. Both approaches were prompted with guidelines. First, we compared the training loss of both approaches. Figure 7 shows that when fine-tuning all the parameters, the loss decreases much more rapidly than when training only the LoRA layers. It also achieves a lower loss at the end of training. However, when evaluating the model at the end of the first and third epochs, we observed that training the full model performs very poorly, as shown in Table 11. We hypothesize that when training all the parameters, the model overfits quickly (indicated by the lower training loss) and memorizes the training data. On the other hand, training only the LoRA layers, which represent around 0.5% of the total model weights, introduces a bottleneck that prevents the model from memorizing the training dataset. It is also noteworthy that the QLoRA approach was trained using just one Nvidia A100 80GB GPU thanks to 4 bits quantization of the frozen model Dettmers et al. (2023). Training the full model required a minimum of four Nvidia A100 80GB GPUs to fit the model into memory. We used DeepSpeed 333github.com/microsoft/DeepSpeed to distribute the model across the four GPUs for training. Due to the high cost of training, we did not perform an extensive hyper-parameter search for the full model.
@@ Appendix E Handling datasets with hundreds of labels and Code-style prompt overhead
Our approach uses Python based code-style prompts, this requires to include tokens in the input to represent the code structures. Figure 8 illustrates various labels formatted in our code-style input alongside their respective guidelines. For very generic guidelines, such as those for the PERSON entity in OntoNotes 5, the code structure accounts for almost half of the input’s characters. However, for detailed guidelines, like those for the Point entity in HarveyNER, the code structure constitutes only a small portion of the input. While there is an overhead of tokens to represent the code structure, when dealing with datasets with a very large number of labels, the primary limitation is fitting the guideline definitions into the model’s input, rather than accommodating the Python code structure.
@@ Appendix G Data-contamination statement
We believe data-contamination is a relevant problem that affects the NLP evaluations nowadays, becoming more prevalent with LLMs (Dodge et al., 2021; Magar & Schwartz, 2022; Sainz et al., 2023b; a). Detecting whether a dataset was inside a LLM pretrained corpora is challenging even with the pre-training data itself. In this paper, unfortunately, we do not have access to the pre-training data used to train Code-LLaMA the backbone LLM of our model. This issue is particularly worrying for us because one big source of contamination is probably GitHub and other code repositories which are also used to upload evaluation benchmarks Dodge et al. (2021). As Code-LLaMA is trained on code, there is a chance for this particular data-leakage. However, all of our comparisons were made against our baseline, which has the same backbone LLM as GoLLIE. Even if the results were impacted, the improvements of our model over the baseline would not be affected by data-contamination as both share the same pre-training. We hope that in the future more transparency will allow to perform safer evaluations.

### 06 Mar 2024 v5

@@ 5.3 Ablation study
We have performed an ablation to see the contribution of several components in the zero-shot evaluation. We analyzed the different regularization techniques proposed in Section 3.3. Additionally, we represent the baseline, i.e. when removing all components including guidelines, as ”w/o all”.
@@
○ GoLLIE.
The class order shuffling, guideline paraphrasing, and class name masking seem to have no significant contribution to the final result, while class dropout although significant improvements are small. As further explained in Appendix D, the loss is only computed over the result tokens, inherently limiting the model’s potential to overfit to the guidelines. In contrast, the representative annotation items give a stronger signal to the model. We see how definitions and representative candidates from the guidelines are complementary and help to improve each other.
@@ 6 Error analysis
The CrossNER dataset introduces two labels for person names within each domain. For example, in the Science domain, the labels Scientist and Person are used. The former is used to label any person that is not a Scientist. Similarly, the Literature domain includes the labels Writer and Person. The guidelines assist GoLLIE in correctly labeling entities as Writer. However, GoLLIE still categorizes individuals as Person even when they are Scientist, despite the guidelines. This is not technically incorrect, as every scientist is, by definition, also a person.
@@
In its Political domain set, CrossNER includes the label Political Party. GoLLIE outperforms the baseline, once again demonstrating the utility of providing the model with guidelines. However, we often find that the model categorizes political parties as organizations. As listed in Table 1, most of the pre-training datasets originate from the news domain, where political parties are a common entity. However, none of the fine-tuning datasets include the Political Party entity; they are instead categorized as Organization. Consequently, during inference, the model consistently labels political parties as organizations. We believe this issue can be resolved by expanding the number and diversity of the fine-tuning datasets.
In summary, we anticipate that GoLLIE will perform well on labels with well-defined and clearly bounded guidelines. On the other hand, ambiguous labels or very coarse labels pose challenges. In this regard, we believe that GoLLIE would benefit from learning to follow instructions such as ”Label always the most specific class” or ”Annotate this class in the absence of other specific class”. We also expect that GoLLIE would benefit from expanding the number and diversity of the pre-training datasets.
@@ 7 Conclusions
○ GoLLIE, an LLM specifically fine-tuned to comply with annotation guidelines that were devised to help humans annotate the dataset. A comprehensive zero-shot evaluation empirically demonstrates that annotation guidelines are of great value for LLMs, as GoLLIE successfully leverages them. GoLLIE achieves better zero-shot results than previous attempts at zero-shot IE which do not leverage the guidelines, or use models not finetuned for following guidelines.
@@ Acknowledgments
This work has been partially supported by the Basque Government (Research group funding IT-1805-22 and ICL4LANG project, grant no. KK-2023/00094). We are also thankful to several MCIN/AEI/10.13039/501100011033 projects: (i) DeepKnowledge (PID2021-127777OB-C21) and by FEDER, EU; (ii) Disargue (TED2021-130810B-C21) and European Union NextGenerationEU/PRTR; (iii) AWARE (TED2021-131617B-I00) and European Union NextGenerationEU/PRTR. This work has also been partially funded by the LUMINOUS project (HORIZON-CL4-2023-HUMAN-01-21-101135724). Oscar Sainz is supported by a doctoral grant from the Basque Government (PRE_2023_2_0137). Rodrigo Agerri currently holds the RYC-2017-23647 fellowship (MCIN/AEI/10.13039/501100011033 and by ESF Investing in your future).
@@ Appendix A Examples
In addition to NER, EE and EAE for which examples are shown in Figures 2 and 3 respectively, we also feed the model with data from RE and SF. The formulation of RE is similar to the NER but with two argument attributes. However, the SF task is more complex as shown in Figure 6. With this task, we added several layers of complexity to the input: extended definitions for each possible attribute (slot), optional arguments, and, fine-grained definitions of types such as Names, Values, or Strings. We also added constraints into the prompt to condition the model to just output the information of our desired query instead of every single template on the text. In the future, we would like to add more complex tasks to the training and evaluation to improve the capabilities and flexibility of the model. For more examples refer to the GitHub repository.
@@ A.1 Example of generalization to new custom Tasks
Our model allows the user to define custom annotation schemas using Python code. We provide an example where we define two new types of entities: Launcher and Mission. As shown in Figure 5, Launcher and Mission are not simple entities, they correspond to what we call Template, a class similar to Entity but with additional arguments, like the SF task. For example, the space_company or the crew of the launcher are some of the additional arguments we added to the schema. As shown in the example, the model’s output (everything after result = [) satisfies the type constraints defined in the guidelines, attributes defined as strings are filled with strings and, the arguments defined as lists (like crew) are filled with lists. The model can correctly analyze the given sentence with our newly created annotation schema.
@@ D.1 Loss calculation
We have used the standard Next Token Prediction (NTP) loss to train our models. However, several regularizations that we applied to the models made the loss computed over the guideline tokens much higher than the actual output tokens. This is because we randomly shuffle the guidelines order, mask names, or, drop classes, which makes it impossible to predict what goes next. To avoid the loss of the guideline tokens overshadowing the actual output tokens loss, we decided to only compute the loss over the output tokens. This way, we can also avoid some overfitting of the guidelines. This resulted in faster training and better results overall.
@@ D.2 Dataset Detailts
Table 9 shows the number of examples for each training and zero-shot dataset. OntoNotes is generated semi-automatically and is orders of magnitude larger than the other ones. Therefore, for each training epoch, we sample 30.000 random examples from the training set. The models were trained for 3 epochs with an effective batch size of 32 and a learning rate of 3e-4 with a cosine scheduler. Therefore, we perform 15,485 training steps.
Regarding the splits, we use the standard train, dev test splits for every dataset. In the case of ACE, we follow the split provided by Lin et al. (2020). In the case of CASIE, we took the first 200 instances as validation and the last 2000 as test.
@@ D.4 LoRa vs Full Model Fine-Tuning
We conducted preliminary experiments to compare the performance of QLoRA (Hu et al., 2022; Dettmers et al., 2023) with that of training all the parameters in the model. These preliminary experiments were conducted using the LLaMA2 7B model Touvron et al. (2023b) and an early version of the code. However, the experimental setup for both models was identical. Both approaches were prompted with guidelines. First, we compared the training loss of both approaches. Figure 7 shows that when fine-tuning all the parameters, the loss decreases much more rapidly than when training only the LoRA layers. It also achieves a lower loss at the end of training. However, when evaluating the model at the end of the first and third epochs, we observed that training the full model performs very poorly, as shown in Table 11. We hypothesize that when training all the parameters, the model overfits quickly (indicated by the lower training loss) and memorizes the training data. On the other hand, training only the LoRA layers, which represent around 0.5% of the total model weights, introduces a bottleneck that prevents the model from memorizing the training dataset. It is also noteworthy that the QLoRA approach was trained using just one Nvidia A100 80GB GPU thanks to the 4-bit quantization of the frozen model Dettmers et al. (2023). Training the full model required a minimum of four Nvidia A100 80GB GPUs to fit the model into memory. We used DeepSpeed 333github.com/microsoft/DeepSpeed to distribute the model across the four GPUs for training. Due to the high cost of training, we did not perform an extensive hyperparameter search for the full model.
@@ Appendix E Handling datasets with hundreds of labels and Code-style prompt overhead
Our approach uses Python-based code-style prompts, this requires including tokens in the input to represent the code structures. Figure 8 illustrates various labels formatted in our code-style input alongside their respective guidelines. For very generic guidelines, such as those for the PERSON entity in OntoNotes 5, the code structure accounts for almost half of the input’s characters. However, for detailed guidelines, like those for the Point entity in HarveyNER, the code structure constitutes only a small portion of the input. While there is an overhead of tokens to represent the code structure, when dealing with datasets with a very large number of labels, the primary limitation is fitting the guideline definitions into the model’s input, rather than accommodating the Python code structure.
@@ Appendix G Data-contamination statement
We believe data contamination is a relevant problem that affects the NLP evaluations nowadays, becoming more prevalent with LLMs (Dodge et al., 2021; Magar & Schwartz, 2022; Sainz et al., 2023b; a). Detecting whether a dataset was inside an LLM pre-trained corpora is challenging even with the pre-training data itself. In this paper, unfortunately, we do not have access to the pre-training data used to train Code-LLaMA the backbone LLM of our model. This issue is particularly worrying for us because one big source of contamination is probably GitHub and other code repositories which are also used to upload evaluation benchmarks Dodge et al. (2021). As Code-LLaMA is trained on code, there is a chance for this particular data leakage. However, all of our comparisons were made against our baseline, which has the same backbone LLM as GoLLIE. Even if the results were impacted, the improvements of our model over the baseline would not be affected by data contamination as both share the same pre-training. We hope that in the future more transparency will allow to perform safer evaluations.